OpenStudy (anonymous):

Solve the following equation by completing the square. x^2+8x+7=0 Find x

sam (.sam.):

x^2+8x+7=0 \[(x+4)^2-16-7=0\] \[(x+4)^2-23=0\] \[(x+4)^2=23\] \[x+4= \pm \sqrt{23}\] Subtract both sides by 4 and you're done

OpenStudy (anonymous):

x^2+8x+7=0 x^2+2*x*4+16-16+7=0 [i have added 16 and subtracted 16 not altering the eqn] (x^2+2*x*4+16)-16+7=0 (x+4)^2 -9=0 (x+4)^2 =9=(+-3)^2 x+4=+-3 x=(+-3)-4 x=(+3-4) or (-3-4) x=-1 or -7

OpenStudy (anonymous):

Got it?

OpenStudy (anonymous):

yes, thank you, I did get it :)